Find the present value. Round to the nearest cent.
To get $10,000 after 2 years at 18% compounded monthly.
Possible answers:
a. $5000.00
b. $6995.44
c. $8363.87
d. $11,956.18

Answer by rfer(16322)

P=A/(1+p/n)^nt
P=10000/(1+0.18/12)^12*2
P=10000/(1.015)^24
P=10000/1.4295
P=$6995.44 (b)